# Strict Optimality of Frequency Estimation Under Local Differential Privacy

### 12 Mar 2026 v1

@@ 4 Communication Cost
This section evaluates the upper bound of the communication cost for an optimal frequency estimator. Let $\mathcal{O}$ represent all potential responses from a LDP client, and $p_{o}$ denote the base probability of a given response $o$. To prevent privacy leakage through the encoding process, we require all responses to utilize the same encoding scheme, regardless of the original value that generated them. Therefore, responses are encoded based on their prior distribution. Shannon’s Source Coding Theorem [shannon1948mathematical] establishes that if the prior distribution of $\mathcal{O}$ is known as $\rho_{o}$ for any response $o$, the optimal communication cost for $\mathcal{O}$ is $\sum_{o}\rho_{o}\log_{2}(\rho_{o})$. In scenarios where the prior distribution is unknown, the upper bound of the optimal communication cost is $\log_{2}(|\mathcal{O}|)$, which occurs when $\forall o:\rho_{o}=\frac{1}{|\mathcal{O}|}$, i.e, logarithm of the number of possible responses.
@@
The communication cost of an optimal frequency estimator can be as low as $\log_{2}(\frac{d(d-1)}{2}+1)$.
The detailed proof is provided in Appendix C. The core idea is to demonstrate that at most $\frac{d(d-1)}{2}+1$ responses are necessary to construct an optimal frequency estimator with an optimal symmetric configuration (Definition 3.1). To derive this bound, we prove that satisfying a symmetric configuration requires that $\forall i<j:\sum_{o}S_{o,i}S_{o,j}p_{o}=\frac{k(k-1)}{(d-1)[k(e^{\epsilon}-1)+d]}$ holds true. This corresponds to $\frac{d(d-1)}{2}$ distinct constraints. By leveraging Carathéodory’s theorem [barvinok2025course], we establish that only $\frac{d(d-1)}{2}+1$ responses are needed; thus, the upper bound of the communication cost is $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$.
There are special cases where the communication cost can be further reduced to $\log_{2}d$. For example, when $d=2^{i}$ for $i\in\mathbb{N}^{+}$ and $k=\frac{d}{2}$ (i.e., as $\epsilon\rightarrow 0$), the support matrix $S$ can be constructed as a Hadamard matrix with $d$ rows, each maintaining a constant support size of $\frac{d}{2}$. This construction is equivalent to the Hadamard Response [acharya2019hr]. Furthermore, assuming the base probability $p_{o}$ is constant across all responses, the construction of SS is equivalent to a Balanced Incomplete Block Design (BIBD). For more details, refer to [bose1939construction]. Depending on the specific values of $(d,k)$, there may exist BIBDs that require only $d$ responses.
Although it is always preferable to identify the optimal communication cost, i.e., a tight lower bound, which this paper does not derive, a communication cost of $O(\log d)$ remains practically feasible even for a very large dictionary size $d$.
@@ 5.1 Subset Selection
This mechanism naturally follows an optimal symmetric configuration. Consequently, by applying the optimized $k$ from Theorem 3.4, Subset Selection serves as an optimal frequency estimator. Given that it possesses $\binom{d}{k}$ possible responses with identical base probabilities, the communication cost is bounded by $k\log_{2}\left(\frac{ed}{k}\right)$ bits, following the inequality $\binom{d}{k}\leq\left(\frac{ed}{k}\right)^{k}$.
@@ 5.2 Optimized Count Mean Sketch
    Determine the hash range $m$ (the number of possible hashed values) based on the privacy parameter $\epsilon$.
@@
    Uniformly sample a hash function $h$ from a hash family $\mathcal{H}$, and map the original value $x$ of an object to a hashed value $y=h(x)$.
@@
    Apply a randomized response mechanism to perturb $y$ within the range $\{1,2,\dots,m\}$, generating a perturbed value $z$. The tuple $(h,z)$ is then sent to the server.
@@
    Frequencies are estimated based on the pairs $(h,z)$ collected from all objects, using a formula equivalent to Eq. (5) as provided by [pan2025improving].
This section demonstrates that CMS is a nearly optimal frequency estimator, subject to the following minor modifications: (1) expanding the original dictionary size $d$ to $d^{\prime}$, where $d^{\prime}$ is the smallest prime number such that $d^{\prime}\geq d$; (2) set hash range $m=\text{round}(1+e^{\epsilon})$ and (3) constructing the hash family as:
$\mathcal{H}=\{h(x)=((ax+b)\mod{d^{\prime}})\mod m|\\ a\in[1,d^{\prime}-1],b\in[0,d^{\prime}-1]\}.$
This construction differs from [pan2025improving] by explicitly disallowing $a=0$. This modification improves the collision probability from approximately $\frac{1}{m^{2}}$ to approximately $\frac{d^{\prime}-m}{m(d^{\prime}-1)}$ (see Appendix D for details). We refer to this variant as Optimized Count Mean Sketch (OCMS).
@@
The output distribution of OCMS with dictionary size $d^{\prime}$ and $m$ buckets is equivalent to sampling between two optimal symmetric configurations:
  • A configuration with support size $\lfloor\frac{d^{\prime}}{m}\rfloor+1$, sampled with probability $p_{\alpha}=\frac{d^{\prime}\mod m}{m}$, and
  • A configuration with support size $\lfloor\frac{d^{\prime}}{m}\rfloor$, sampled with probability $1-p_{\alpha}$.
The detailed proof is provided in Appendix D. The core intuition is that any response $(h,z)$ supports all elements mapped to bucket $z$. Each bucket contains either $\lfloor\frac{d^{\prime}}{m}\rfloor+1$ or $\lfloor\frac{d^{\prime}}{m}\rfloor$ elements. By grouping all possible responses according to their support size, each group constitutes a distinct configuration; thus, OCMS can be reconstructed as a weighted sampling of these two configurations.
This lead to the following result:
@@
$\text{Var}(\hat{f}_{OCMS}(x))=p_{\alpha}Var(\hat{f}^{*}(x|k=\lfloor\frac{d^{\prime}}{m}\rfloor+1))+\\ (1-p_{\alpha})Var(\hat{f}^{*}(x|k=\lfloor\frac{d^{\prime}}{m}\rfloor)),$
@@
Define $\alpha=\frac{(d-1)(e^{\epsilon}+1)^{4}}{(4de^{\epsilon}-(e^{\epsilon}+1)^{2})(d-e^{\epsilon}-1)(de^{\epsilon}+e^{\epsilon}+1)}$, $\beta=\frac{(d^{\prime}-d)(e^{\epsilon}+1)^{2}(2dd^{\prime}-d-d^{\prime})}{d^{\prime 2}(d-1)(4de^{\epsilon}-(e^{\epsilon}+1)^{2})}$. If OCMS has an optimized hash range $m=\text{round}(1+e^{\epsilon})$ , then
$\mathcal{L}_{1,OCMS}^{*}\leq\mathcal{L}_{1}^{*}\sqrt{(1+\alpha)(1+\beta)}$
$\mathcal{L}_{2,OCMS}^{*}\leq\mathcal{L}_{2}^{*}(1+\alpha)(1+\beta)$
where $\mathcal{L}_{1}^{*}$ and $\mathcal{L}_{2}^{*}$ represent the optimized $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ given dictionary size $d$ and privacy factor $\epsilon$, respectively.
The proof is detailed in Appendix D. Notably, when $m=\text{round}(1+e^{\epsilon})$, OCMS is equivalent to sampling from two optimal symmetric configurations with support sizes of $\lfloor\frac{d^{\prime}}{1+e^{\epsilon}}\rfloor$ and $\lceil\frac{d^{\prime}}{1+e^{\epsilon}}\rceil$, respectively. In the regime where $d\gg e^{\epsilon}$, both $\alpha$ and $\beta$ vanish toward zero. For instance, with $\epsilon=1$, $d=100$, and $d^{\prime}=101$, the $\mathcal{L}_{2}$ error of OCMS is only $0.09\%$ higher than the theoretical lower bound.
@@ 5.3 Weighted Subset Selection
Theorem 4.1 proves that at most $\frac{d(d-1)}{2}+1$ responses are required to construct an optimized frequency estimator. These responses constitute a subset of $S^{\star}$, which denotes the set of all possible $\binom{d}{k}$ combinations. We can develop an algorithm to identify these $\frac{d(d-1)}{2}+1$ responses and construct the corresponding matrices SS and $P$. The core construction procedure is as follows:
@@
In our experiments, this heuristic typically generates a valid $S$ when $m=d^{2}$. The constraint $A\mathbf{p}=y$ inherently guarantees an optimal symmetric configuration (see the proof of Theorem 4.1). Consequently, it serves as an optimized frequency estimator provided that the support size $k$ is optimized. Since LP or NNLS solvers tend to produce sparse solutions, we observe that $\mathbf{p}$ generally contains no more than $\frac{d(d-1)}{2}+1$ positive values. Thus, the communication cost is typically bounded by $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$. If a specific solution for $\mathbf{p}$ contains more responses than this bound, the process can be restarted to find a sparser alternative.
Notably, LP or NNLS solvers have a computational complexity of $O(H^{2}W)$, where $H$ and $W$ represent the height and width of $A$, respectively. In this context, $H=\frac{d(d-1)}{2}$ and $W=m=O(d^{2})$, resulting in a total computational complexity of $O(d^{6})$. A practical approach to adopting WSS is to precompute valid support matrices $S$ for common values of $d$ and $k$; clients can then simply download and reuse these matrices for each round of frequency estimation. Furthermore, as the proof of Corollary 5.2 demonstrates that extending a dictionary to a larger size has a negligible impact on precision, one can simply use the next available precomputed $d$ for which a valid $S$ exists.
@@ 5.4 Choosing an Optimal Algorithm
When selecting an optimal frequency estimator, the choice largely depends on the dictionary size $d$. If $d$ is small, WSS or SS are viable options; otherwise, OCMS is preferable due to its low communication and precomputation overhead. To determine whether a dictionary is sufficiently large for OCMS to be effective, one may refer to Corollary 5.2. For example, when $\epsilon=1$ and $d\geq 100$, $\mathcal{L}_{2}$ of OCMS is at most 0.09% above the optimal loss, a gap that continues to narrow as $d$ increases.
Additionally, if the parameters of WSS have been precomputed for a specific privacy factor $\epsilon$ and a dictionary size larger than your own, you may extend your dictionary to leverage these precomputed values.
Figure 1: $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses vs. privacy factor $\epsilon$ given the Zipf dataset. See Section 6.1 for details.
Figure 2: $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses vs. privacy factor $\epsilon$ given the mini Kosarak dataset. See Section 6.2 for details.
@@ 6 Experiment
This section empirically evaluates if the aforementioned optimal algorithms achieve the theoretical $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses.
We conduct two experiments, each consisting of 100 independent runs. The $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses of the evaluated algorithms are measured as:
$L_{1}=\frac{1}{t}\sum_{t}\sum_{x\in[d]}|\hat{f}(x)-f(x)|$
$L_{2}=\frac{1}{t}\sum_{t}\sum_{x\in[d]}(\hat{f}(x)-f(x))^{2}$
where $t$ represents the number of experiment runs. Since $\mathbb{E}[L_{1}]=\mathcal{L}_{1}$ and $\mathbb{E}[L_{2}]=\mathcal{L}_{2}$, these measures are unbiased and should closely approximate the theoretical $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ when $t$ is large enough.
6.1 Zipf’s Distribution
The first experiment utilizes a synthetic dataset sampled from a specific Zipf’s distribution, where the frequency distribution is defined as:
$\forall x\in[d]:Zipf(x)\propto\frac{1}{x^{2}},$
@@ 6.2 Real-world Dataset
The second experiment is conducted on Kosarak [kosarak], a real-world dataset containing click-stream data from a Hungarian online news portal. For our experiment, we reduce this dataset to 1% of its original size (denoted as "mini Kosarak"), which still contains over 170,000 objects with a dictionary size $d=26,000$. In this setup, we only evaluate Subset Selection (SS) and Optimized Count Mean Sketch (OCMS). Weighted Subset Selection (WSS) is omitted due to its high precomputation cost for a dictionary of this size. The results are plotted in Fig. 2.

### 22 Mar 2026 v2

@@ 4 Communication Cost
This section evaluates the upper bound of the communication cost for an optimal frequency or distribution estimator. Let $\mathcal{O}$ represent all potential responses from a LDP client, and $p_{o}$ denote the base probability of a given response $o\in\mathcal{O}$. To prevent privacy leakage through the encoding process, we require all responses to utilize the same encoding scheme, regardless of the original value that generated them. Therefore, responses are encoded based on their prior distribution. Shannon’s Source Coding Theorem [shannon1948mathematical] establishes that if the prior distribution of $\mathcal{O}$ is known for any response $o$ as $\rho_{o}$, the optimal communication cost for $\mathcal{O}$ is $\sum_{o}\rho_{o}\log_{2}(\rho_{o})$ bits. In scenarios where the prior distribution is unknown, the upper bound of the optimal communication cost is $\log_{2}(|\mathcal{O}|)$ bits, which occurs when $\forall o:\rho_{o}=\frac{1}{|\mathcal{O}|}$, i.e., the logarithm of the number of possible responses.
@@
The communication cost of an optimal estimator can be as low as $\log_{2}(\frac{d(d-1)}{2}+1)$ bits.
The detailed proof is provided in Appendix C. The core idea is to demonstrate that at most $\frac{d(d-1)}{2}+1$ responses are necessary to construct an optimal estimator with an optimal symmetric configuration (Definition 3.2). To derive this bound, we prove that satisfying a symmetric configuration requires that $\forall i<j:\sum_{o}S_{o,i}S_{o,j}p_{o}=\frac{k(k-1)}{(d-1)[k(e^{\epsilon}-1)+d]}$ holds true. This corresponds to $\frac{d(d-1)}{2}$ distinct constraints. By leveraging Carathéodory’s theorem [barvinok2025course], we establish that only $\frac{d(d-1)}{2}+1$ responses are needed; thus, the upper bound of the communication cost is $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$ bits.
There are special cases where the communication cost can be further reduced to $\log_{2}d$. For example, when $d=2^{i}$ for $i\in\mathbb{N}^{+}$ and $k=\frac{d}{2}$ (i.e., as $\epsilon\rightarrow 0$), the support matrix $S$ can be constructed as a Hadamard matrix with $d$ rows, each maintaining a constant support size of $\frac{d}{2}$. This construction is equivalent to the Hadamard Response [acharya2019hr]. Furthermore, assuming the base probability $p_{o}$ is constant across all responses, the construction of SS is equivalent to a Balanced Incomplete Block Design (BIBD). For more details, refer to [bose1939construction]. Depending on the specific values of $(d,k)$, there may exist BIBDs that require only $d$ responses [park2024exactly].
Although it is always preferable to identify the optimal communication cost, i.e., a tight lower bound, which this paper does not derive, a communication cost of $O(\log d)$ bits remains practically feasible even for a very large dictionary size $d$.
@@ 5.1 Subset Selection
This mechanism naturally follows an optimal symmetric configuration. Consequently, by applying the optimized $k$ from Theorem 3.8, Subset Selection serves as an optimal estimator. Given that it possesses $\binom{d}{k}$ possible responses with identical base probabilities, the communication cost is bounded by $k\log_{2}\left(\frac{ed}{k}\right)$ bits, following the inequality $\binom{d}{k}\leq\left(\frac{ed}{k}\right)^{k}$.
@@ 5.2 Optimized Count Mean Sketch
    Determine the hash range $B$ (the number of possible hashed values) based on the privacy parameter $\epsilon$.
@@
    Uniformly sample a hash function $h$ from a hash family $\mathcal{H}$, and map an object’s original value $x$ to a hashed value $y=h(x)$.
@@
    Apply a randomized response mechanism to perturb $y$ within the range $\{1,2,\dots,B\}$, generating a perturbed value $z$. The tuple $(h,z)$ is then sent to the server.
@@
    Frequencies or distributions are estimated based on the pairs $(h,z)$ collected from all objects, using a formula equivalent to Eq. (5) as provided by [pan2025improving].
This section demonstrates that CMS is a nearly optimal frequency or distribution estimator, subject to the following minor modifications: (1) expanding the original dictionary size $d$ to $d^{\prime}$, where $d^{\prime}$ is the smallest prime number such that $d^{\prime}\geq d$; (2) set hash range $m=\text{round}(1+e^{\epsilon})$ and (3) constructing the hash family as:
$\mathcal{H}=\{h(x)=((ax+b)\mod{d^{\prime}})\mod B|\\ a\in[1,d^{\prime}-1],b\in[0,d^{\prime}-1]\}.$
This construction differs from [pan2025improving] by explicitly disallowing $a=0$. This modification improves the collision probability from approximately $\frac{1}{B^{2}}$ to approximately $\frac{d^{\prime}-B}{B^{2}(d^{\prime}-1)}$ (see Appendix D.1 for details). We refer to this variant as Optimized Count Mean Sketch (OCMS).
@@
The output distribution of OCMS with dictionary size $d^{\prime}$ and $B$ buckets is equivalent to sampling between two optimal symmetric configurations:
  • A configuration with support size $\lfloor\frac{d^{\prime}}{B}\rfloor+1$, sampled with probability $p_{\alpha}=\frac{d^{\prime}\mod B}{B}$, and
  • A configuration with support size $\lfloor\frac{d^{\prime}}{B}\rfloor$, sampled with probability $1-p_{\alpha}$.
The detailed proof is provided in Appendix D. The core intuition is that any response $(h,z)$ supports all elements mapped to bucket $z$. Each bucket contains either $\lfloor\frac{d^{\prime}}{B}\rfloor+1$ or $\lfloor\frac{d^{\prime}}{B}\rfloor$ elements. By grouping all possible responses according to their support size, each group constitutes a distinct configuration; thus, OCMS can be reconstructed as a weighted sampling of these two configurations.
This leads to the following result:
@@
$\text{Var}(\hat{f}_{OCMS}(x))=p_{\alpha}Var(\hat{f}^{*}(x|k=\lfloor\frac{d^{\prime}}{B}\rfloor+1))+\\ (1-p_{\alpha})Var(\hat{f}^{*}(x|k=\lfloor\frac{d^{\prime}}{B}\rfloor)),$
@@
Define $\alpha=\frac{(d-1)(e^{\epsilon}+1)^{4}}{(4de^{\epsilon}-(e^{\epsilon}+1)^{2})(d-e^{\epsilon}-1)(de^{\epsilon}+e^{\epsilon}+1)}$, $\beta=\frac{(d^{\prime}-d)(e^{\epsilon}+1)^{2}(2dd^{\prime}-d-d^{\prime})}{d^{\prime 2}(d-1)(4de^{\epsilon}-(e^{\epsilon}+1)^{2})}$. If OCMS has an optimized hash range $B=\text{round}(1+e^{\epsilon})$ , then
$\mathcal{L}_{1,OCMS}^{*}(n)\leq\mathcal{L}_{1}^{*}(n)\sqrt{(1+\alpha)(1+\beta)}$
$\mathcal{L}_{2,OCMS}^{*}(n)\leq\mathcal{L}_{2}^{*}(n)(1+\alpha)(1+\beta)$
where $\mathcal{L}_{1}^{*}$ and $\mathcal{L}_{2}^{*}$ represent the optimized $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ given dictionary size $d$ and privacy parameter $\epsilon$, respectively.
The proof is detailed in Appendix D. Notably, when $B=\text{round}(1+e^{\epsilon})$, OCMS is equivalent to sampling from two optimal symmetric configurations with support sizes of $\lfloor\frac{d^{\prime}}{1+e^{\epsilon}}\rfloor$ and $\lceil\frac{d^{\prime}}{1+e^{\epsilon}}\rceil$, respectively. In the regime where $d\gg e^{\epsilon}$, both $\alpha$ and $\beta$ vanish toward zero. For instance, with $\epsilon=1$, $d=100$, and $d^{\prime}=101$, the $\mathcal{L}_{2}$ error of OCMS is only $0.09\%$ higher than the theoretical lower bound.
@@ 5.3 Weighted Subset Selection
Theorem 4.1 proves that at most $\frac{d(d-1)}{2}+1$ responses are required to construct an optimal estimator. These responses constitute a subset of $S^{\star}$, which denotes the set of all possible $\binom{d}{k}$ combinations. We can develop an algorithm to identify these $\frac{d(d-1)}{2}+1$ responses and construct the corresponding matrices SS and $P$. The core construction procedure is as follows:
@@
In our experiments, this heuristic typically generates a valid $S$ when $m=d^{2}$. The constraint $A\mathbf{p}=y$ inherently guarantees an optimal symmetric configuration (see the proof of Theorem 4.1). Consequently, it serves as an optimal estimator provided that the support size $k$ is optimized. Since LP or NNLS solvers tend to produce sparse solutions, we observe that $\mathbf{p}$ generally contains no more than $\frac{d(d-1)}{2}+1$ positive values. Thus, the communication cost is typically bounded by $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$. If a specific solution for $\mathbf{p}$ contains more responses than this bound, the process can be restarted to find a sparser alternative.
Notably, LP or NNLS solvers have a computational complexity of $O(H^{2}W)$, where $H$ and $W$ represent the height and width of $A$, respectively. In this context, $H=\frac{d(d-1)}{2}$ and $W=m=O(d^{2})$, resulting in a total computational complexity of $O(d^{6})$. A practical approach to adopting WSS is to precompute valid support matrices $S$ for common values of $d$ and $k$; clients can then simply download and reuse these matrices for each round of frequency or distribution estimation. Furthermore, as the proof of Corollary 5.2 demonstrates that extending a dictionary to a larger size has a negligible impact on precision, one can simply use the next available precomputed $d$ for which a valid $S$ exists.
@@ 5.4 Choosing an Optimal Algorithm
When selecting an optimal frequency or distribution estimator, the choice largely depends on the dictionary size $d$. If $d$ is small, WSS or SS are viable options; otherwise, OCMS is preferable due to its low communication and precomputation overhead. To determine whether a dictionary is sufficiently large for OCMS to be effective, one may refer to Corollary 5.2. For example, when $\epsilon=1$ and $d\geq 100$, $\mathcal{L}_{2}$ of OCMS is at most 0.09% above the optimal loss, a gap that continues to narrow as $d$ increases.
Additionally, if the parameters of WSS have been precomputed for a specific privacy parameter $\epsilon$ and a dictionary size larger than your own, you may extend your dictionary to leverage these precomputed values.
Figure 1: $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses vs. privacy parameter $\epsilon$ given the Zipf distribution. See Section 6.1 for details.
Figure 2: $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses vs. privacy parameter $\epsilon$ given the mini Kosarak dataset. See Section 6.2 for details.
@@ 6 Experiment
This section empirically evaluates if the aforementioned optimal algorithms achieve the theoretical $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses. We conduct two experiments, each consisting of 100 independent runs. The $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses of the evaluated algorithms are measured as:
$L_{1}=\frac{1}{t}\sum_{t}\sum_{x\in[d]}|\hat{r}(x)-r(x)|$
$L_{2}=\frac{1}{t}\sum_{t}\sum_{x\in[d]}(\hat{r}(x)-r(x))^{2}$
where $t$ represents the number of experiment runs, $r$ can be either frequency $f$ or distribution $\theta$. Since $\mathbb{E}[L_{1}]=\mathcal{L}_{1}$ and $\mathbb{E}[L_{2}]=\mathcal{L}_{2}$, these measures are unbiased and should closely approximate the theoretical $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ when $t$ is large enough.
6.1 Zipf Distribution
The first experiment samples datasets from a specific Zipf distribution to estimate the distribution $\bm{\theta}$, which is defined as:
$\forall x\in[d]:\theta(x)=\text{Zipf}(x)\propto\frac{1}{x^{2}},$
@@ 6.2 Real-world Dataset
The second experiment estimates the frequency of Kosarak [kosarak], a real-world dataset containing click-stream data from a Hungarian online news portal. For our experiment, we reduce this dataset to 1% of its original size (denoted as "mini Kosarak"), which still contains over 170,000 objects with a dictionary size $d=26,000$. In this setup, we only evaluate Subset Selection (SS) and Optimized Count Mean Sketch (OCMS). Weighted Subset Selection (WSS) is omitted due to its high precomputation cost for a dictionary of this size. The results are plotted in Fig. 2.